OpenStudy (anonymous):

choose the equation of the line with slope=-4 and containing the point (-3,-4)

OpenStudy (anonymous):

use this \[\huge y=mx+b\]

OpenStudy (anonymous):

since the slope is given and one point is given thus the line is y+4=-4(x+3) or y=-4x-12-4 or4x+y+16=0